Use the following information to answer questions 14 through 18. Assume Do-Nothling Alternative #1 Alternative #2 $205,000 $45,000 in year 1; First Cost $100,000 $35,000 per year increase of $3,000 per year S5,000 per years $10,000 per year Benefits O&M Costs Salvage Value Useful Life ROR for each of the following years $10,000 $15,000 years 10.85% 14. The NPW of Alternative #1 (necessary to make a decision between the two a) S3,249 $1,930 $543 e) $2,227 f)-$2,311 8) $723 h) $431 ) $3,860 15. The increase in O&M costs that would result in a 10% ROR for Alternative #1 is closest to a) $483 per year b)S517 per year c) $813 per year d) $609 per year e) $407 per year f) $702 per year g)O&M costs need to decrease in order to earn a ROR of 10% h) not enough information to determine 6-The EU AB for Alternative #2 is closest to discussed in class.) (Remember to treat the salvage value using the standard me a) $37,525 b) $43,81 c) $80,939 d) $72,373 e) $44,592 DS54,012 g) $55,323 17. The EUAC fr Alternative #2 is closest to discussed in class.) (Remember to treat the salvage value using the standard #15
Answer #1

14. Ans: $1,930

Explanation:

NPW = -P -A(P/A, i, n) + A(P/A, i, n) + F(P/F, i, n)

         = -100,000 - 5,000(P/A, 10%, 4) + 35,000(P/A, 10%, 4) + 10,000(P/F, 10%, 4)

         = -100,000 - 5,000(3.170) + 35,000(3.170) + 10,000(0.6830)

         = -100,000 - 15,850 + 110,950 + 6,830

         = $1,930